OpenStudy (timaashorty):

Help would be appreciated(: Find all the zeros of of the equation -3x^4+27x^2+1200=0

OpenStudy (timaashorty):

@satellite73

OpenStudy (timaashorty):

oh there you are (:

OpenStudy (anonymous):

start by finding the zeros of \[-3u^2+27u+1200=0\] where \(u=x^2\)

OpenStudy (anonymous):

best bet is to divide all by \(-3\) and start with \[u^2-9u-400=0\]

OpenStudy (anonymous):

by some miracle this one actually factors as \[(u+16)(u-25)=0\] and so \(u=-16\) or \(u=25\) now go back to \(x^2\) and write \[x^2=-16\] or \[x^2=25\]

OpenStudy (timaashorty):

So I would have to write it like this -3x^4+27(-16)^2+1200=0 ? Or x^2= were my finall answers ?

OpenStudy (anonymous):

no your job is to solve for \(x\) not rewrite it you have \[x^2=25\] so \[x=\pm5\] and also \[x^2=-16\] which is not possible with real numbers

OpenStudy (timaashorty):

Oh isee ! Thank you very much for the help! (:

OpenStudy (anonymous):

yw

OpenStudy (timaashorty):

Thank you timo (: